JustSaiyan:

Is it possible for the solution set of a system of linear inequalities to be all real numbers? Explain your reasoning.

JustSaiyan:

@Arcadiouse

JustSaiyan:

@Angle

Angle:

x + 3y = 2 2x + 6y = 4

JustSaiyan:

Thank you.

Angle:

what do you get when you try solving for x and y?

JustSaiyan:

A head ache?

Angle:

LOL yes but I tried making the example easy ;) you gonna try? :3

Arcadiouse:

Yeah lol. M: Mental A: Abuse T: To H: Humans

JustSaiyan:

This is algebra.

Arcadiouse:

Still math.

JustSaiyan:

x + 7y = 9 5x + 9y = 16

Angle:

ewwww what are you doing? x'D

JustSaiyan:

Pretending to know how this works.

JustSaiyan:

But could you answer the actual question?

Angle:

x + 3y = 2 2x + 6y = 4 you can divide by 2 through the entire second equation and get x + 3y = 2 but that's just the first equation

Angle:

I'm saying that you can have a system of equations where it's just the same equation and the same equation = that same equation = solutions are all real numbers

Angle:

since the equation overlaps itself perfectly, all real numbers are the answer

JustSaiyan:

Thank you.

JustSaiyan:

It seems this site is more helpful than PA, OS, and brainly.

Angle:

but do you understand how I explained it?

JustSaiyan:

Yes.

Angle:

these sorts of equations are really easy to confuse with the "no solution" type of system of equations, so just trying to solve for x and y is your best bet of figuring such equations out ^_^